One of my assessment questions was:

"What is 2.3 x 10 in exponential notation?"
My answer was 2.3 x 10(to the 1st power); which equals 23. However,
their correct answer was 2.3 x 10 (to the zero power).
I thought any base to the zero power is 1 (and 2.3 x 1 = 2.3).
Why is their answer correct?

Answer by jim_thompson5910(35256)


You are correct in both statements that
10^1 = 10, so 2.3 x 10 = 2.3 x 10^1
10^0 = 1, so 2.3 x 10^0 = 2.3 x 1 = 2.3

It seems like your teacher made a typo.